A photo of some of the damage to the fiber cable.
BY MAIRE O’NEILL
The Los Alamos Reporter published a story Monday morning about the weekend cellular and internet service in Los Alamos County. Surprisingly, little information was available at that time. Since then members of the community have reached out with concerns about the lack of information available to the public as well as the County’s lack of means to make the public aware of what actually is going on in a situation like this and how soon service will be resumed.
Initially, only one cut of a fiber cable was reported but later Los Alamos National Laboratory internal reports indicated that there were two cuts. An unofficial Lumen report received by the Reporter indicates the following:
A damaged fiber cable in Los Alamos, NM impacted client services.
The affected fiber was repaired, thus restoring impacted services to a stable state.
Lumen Network Operations Center (NOC) was engaged and investigated the issue. Initial investigations indicated a suspected fiber fault was the cause of disruption. As such, Field Operations was engaged to assist with isolation efforts and arrived onsite of the testing facility by 11:49 GMT. Upon arrival, light-level testing was conducted, and measurement data indicated a fiber break approximately 24 miles from the testing location in Los Alamos towards Santa Fe. At approximately 21:00 GMT the Outside Plant Engineer (OSPE) was engaged and dispatched to the suspected point of failure to assist with ongoing isolation efforts.
Repair teams advised that splice cases were full of mud and water which impeded isolation efforts and the OSPE suspected that multiple fault locations were present. The first fiber fault was isolated, and teams began efforts to complete repairs as another team dispatched to traverse the span and isolate the secondary failure.
On December 12, 2022, at 00:32 GMT, Field Operations advised that the first section of damaged fiber cable had been repaired; however, services remained affected. Field Operations then began taking additional fiber measurements from handholes along the fiber span to isolate the failure. As the teams were working to test the fiber strand by strand, isolation efforts were tedious and further delayed by standing water and mud in handholes in Paseo Estrellas. The fiber fault was finally isolated at 07:17 GMT, and Field Operations advised that a rodent chew was affecting the cable. Teams began repair efforts on the cable and at 09:35 GMT, associated alarms cleared and services restored to a stable condition.